# Suppression of Gravitational-Wave Echoes in Diffeomorphism-Invariant Nonlocal Gravity

### v1

@@ Abstract
Searches for late–time gravitational–wave echoes following compact binary mergers are often interpreted as probes of exotic near–horizon or horizonless physics. We show that a well–motivated class of ultraviolet–finite quantum gravity and modified gravity theories, those based on diffeomorphism–invariant, analytic, entire–function nonlocality, generically suppress observable echo signals. The suppression follows from Paley–Wiener bounds associated with the analyticity of the nonlocal regulator, which enforce exponential damping of high–frequency reflection amplitudes and smear sharp effective potentials in the black–hole interior. We demonstrate that for regular black holes with horizons, the standard ingoing boundary condition eliminates the cavity required for echoes, while for regular horizonless compact objects the nonlocal kernel strongly attenuates both inner and photon–sphere reflections. As a result, repeated reflections during the ringdown phase are exponentially suppressed in frequency space and washed out in the time domain. Our results imply that the absence of echoes in current gravitational–wave data is consistent with covariant nonlocal gravity theories and does not, by itself, favor classical horizons over regular or horizonless ultraviolet completions. The analyticity and Paley–Wiener bounds that enforce echo suppression are required to ensure unitarity, ghost freedom, and ultraviolet finiteness of the underlying quantum theory. As a result, the classical ringdown phenomenology already encodes quantum–gravity consistency conditions. We emphasize that the absence of echoes in such theories does not rely on quantum decoherence or stochastic effects, but instead reflects the analytic structure imposed by quantum gravity on the classical limit.
@@ 1 Introduction
The detection of gravitational waves from compact binary mergers has opened a new observational window onto the strong–field regime of gravity. In particular, the post–merger ringdown phase, dominated by quasinormal modes of the remnant compact object, provides a sensitive probe of the near–horizon structure of spacetime. Motivated by this possibility, a number of works have proposed that departures from classical black holes—such as horizonless compact objects, quantum–modified horizons, or reflective surfaces replacing the event horizon—could generate late–time, repeating echoes in the gravitational–wave signal.
Echoes arise when gravitational perturbations are repeatedly trapped between an effective potential barrier near the photon sphere and a partially reflective inner boundary. In classical general relativity, an event horizon enforces purely ingoing boundary conditions, preventing such a cavity from forming. By contrast, many phenomenological echo models assume ad hoc reflective conditions at microscopic distances outside the would–be horizon or at the surface of a compact object. The observational status of echoes remains unsettled, with current data providing no statistically compelling evidence for their existence.
@@
We show that these same analytic and diffeomorphism–invariant properties have direct and robust consequences for gravitational–wave ringdown phenomenology [13, 14, 15, 16, 17]. In particular, the Paley–Wiener bounds enforce an exponential suppression of high–frequency reflection amplitudes and smear any sharp features in the effective radial potential governing perturbations. As a result, the two essential ingredients required for observable echoes—strong inner reflection and a sharply defined resonant cavity—are generically absent.
We analyze both regular black holes possessing true event horizons and regular horizonless compact objects arising in nonlocal gravity. In the former case, the presence of a horizon enforces ingoing boundary conditions that eliminate inner reflection altogether. In the latter case, regularity at the center replaces the notion of a hard reflecting surface, while the nonlocal kernel suppresses scattering from interior structure and from the photon–sphere barrier itself. In both situations, any would–be echo train is exponentially damped with increasing frequency and echo number, and is further smeared in the time domain by the intrinsic nonlocal scale.
Our results demonstrate that the absence of gravitational–wave echoes is a prediction of analytic entire–function nonlocal gravity theories. Consequently, non–detection of echoes should not be interpreted as evidence against ultraviolet completions featuring regular black holes or horizonless compact objects, but rather as a consistency check of covariant nonlocal dynamics. Ringdown observations instead constrain the nonlocal scale and the analytic structure of the gravitational form factor, providing a complementary probe of quantum gravity in the strong–field regime.
@@
The suppression of gravitational–wave echoes derived in this work operates entirely at the level of classical wave propagation on a fixed background geometry. The modified Regge–Wheeler–type equations [18, 19, 20], the smearing of effective potentials, and the exponential damping of reflection coefficients follow from classical integro–differential equations of motion. However, the specific nonlocal structure responsible for these effects, namely, analytic entire functions of the covariant d’Alembertian obeying Paley–Wiener bounds—is not motivated by classical considerations alone. Such analyticity conditions are imposed to ensure unitarity, ghost freedom, and ultraviolet finiteness of the quantum gravitational theory. In this sense, quantum gravity is already implicitly encoded in the classical dynamics. The same analytic properties required for a consistent quantum theory manifest classically as a frequency–space filter that suppresses sharp reflections and eliminates the resonant cavity necessary for observable echoes. The resulting ringdown phenomenology reflects quantum–gravity consistency conditions, even in the absence of explicit quantum fluctuations or loop effects.
@@ 2 Regular Static Spherically Symmetric Solution
normalized so that the total mass is $M$:
@@
The relation between $\ell$ and $\Lambda_{G}$ depends on convention $\ell=1/\Lambda_{G}$ or $\ell=1/(\sqrt{2}\Lambda_{G}$; for the horizon phase structure only the ratio $GM/\ell$ matters.
@@
$x:=\frac{r}{\ell},\qquad\mu:=\frac{GM}{\ell},,$ (14)
@@
3 Suppression of Ringdown Echoes in Covariant Nonlocal Gravity
@@
where $K_{\Lambda_{G}}$ is a smooth kernel of width $\Delta r_{*}\sim\Lambda_{G}^{-1}$ in the local proper measure, determined by $\mathcal{F}$. Equation (23) states that the nonlocality smears the potential and modifies scattering.
Defining the smeared effective potential:
@@
one may formally rewrite (23) as
$\Big[\partial_{r_{*}}^{2}+\omega^{2}\Big]\Psi_{\omega}(r_{*})\;=\;V_{\rm eff}(r_{*})\,\Psi_{\omega}(r_{*})\;+\;\Delta_{\rm NL}[\Psi_{\omega}],$ (25)
where $\Delta_{\rm NL}$ collects non-multiplicative nonlocal terms; in the Born scattering regime the convolution form (23) is sufficient to establish robust suppression bounds.
Let the Fourier transform of the kernel be given by
$\widetilde{K}_{\Lambda_{G}(k)}\;\equiv\;\int_{-\infty}^{+\infty}dx\;e^{-ikx}K_{\Lambda_{G}}(x).$ (26)
For the Gaussian regulator $n=1$ in (20):
$\widetilde{K}_{\Lambda_{G}(k)}\;=\;\exp\!\Big[-k^{2}/\Lambda_{G}^{2}\Big],$ (27)
and for general $n$ one obtains a function that decays super–polynomially, indeed faster than any power at large $|k|$. More generally, the analyticity and exponential–type bounds on $\mathcal{F}$ imply Paley–Wiener bounds on $\widetilde{K}_{\Lambda_{G}(k)}$ [11, 12], yielding strong damping in the frequency domain and smooth kernels in the radial domain. Physically, the nonlocality acts as a band–limiting filter for scattering.
@@
The analytic nonlocality multiplicatively suppresses reflection through the factor $\widetilde{K}_{\Lambda_{G}}(2\omega)$. For the Gaussian example (27):
@@
$R_{\rm ph,NL}(\omega)\;\approx\;\widetilde{K}_{\Lambda_{G}}(2\omega)\,R_{\rm ph,loc}(\omega),\qquad T_{\rm ph,NL}(\omega)\;\approx\;T_{\rm ph,loc}(\omega)\ \ \text{(to leading order)}.$ (31)
A similar filtering applies to any effective interior barrier that would otherwise reflect waves back toward the photon sphere.
@@
where $R_{\rm in}(\omega)$ is the inner reflection amplitude and $\Delta t$ is the round–trip time between the photon sphere and the inner scattering region expressed in tortoise coordinate. The transfer function is given by
@@
Combining (31) and (37) in the echo series (32), the $n$th echo carries a factor:
@@
$\boxed{\Big|R_{\rm ph,NL}(\omega)\,R_{\rm in,NL}(\omega)\Big|^{n}\;\lesssim\;\exp\!\Big[-\,8n\,\omega^{2}/\Lambda_{G}^{2}\Big]\;\Big|R_{\rm ph,loc}(\omega)\,R_{\rm in,loc}(\omega)\Big|^{n}}$ (40)

### v2

# Suppression of Gravitational-Wave Echoes in Diffeomorphism-Invariant Nonlocal Quantum Gravity

@@ Abstract
Searches for gravitational-wave echoes have been widely interpreted as probes of near-horizon structure and quantum modifications of black holes. We revisit the mechanism by which echoes are suppressed in a diffeomorphism-invariant, analytic entire-function ultraviolet completion of quantum gravity. We show that the absence of observable echoes is not due to a suppression of quasinormal-mode frequencies or a filtering of the gravitational-wave spectrum. The extreme blueshift of the proper local frequency in the near-horizon region activates the diffeomorphism-invariant entire-function regulator, which smooths out sharp reflecting inner structures and drives the reflection coefficient to zero. The nonlocal regulator acts as a covariant smearing operator on effective stress–energy and curvature, replacing sharp, partially reflective surfaces by smooth transition regions. As a result, the would-be echo cavity required for repeated reflections fails to form, and the associated phase-coherent back scattering is eliminated. This mechanism applies both to regular black holes with horizons and to horizonless compact objects generated by nonlocal smearing, and is independent of the detailed value of the nonlocal scale provided it lies at or above the Planck scale. Our results clarify the physical origin of echo suppression in nonlocal quantum gravity and demonstrate that the absence of echoes is a structural consequence of analytic nonlocality rather than a dynamical damping of gravitational-wave frequencies.
@@ 1 Introduction
Gravitational-wave observations of compact binary mergers provide a new observational window on the strong-field regime of gravity. In particular, the post-merger ringdown phase has been proposed as a sensitive probe of near-horizon or near-surface modifications of black holes, including quantum-gravity-inspired deviations from classical general relativity. A prominent phenomenological signature is the appearance of late-time gravitational-wave echoes, generated by partial reflection of perturbations between an effective potential barrier and an inner reflective structure. The absence of statistically significant echoes in current observations has therefore been used to constrain a wide class of exotic compact object models.
@@
Nonlocal, analytic entire-function ultraviolet completions of quantum gravity constitute a distinct theoretical framework in which both ultraviolet finiteness and singularity resolution can be achieved without introducing new degrees of freedom or violating diffeomorphism invariance. In this approach, nonlocality enters through covariant entire functions of the d’Alembertian operator, which act as smoothing operators on stress–energy distributions and curvature invariants. In static and dynamical settings, this smearing replaces otherwise sharp geometric or material features by smooth transition regions, while preserving the classical description of spacetime at macroscopic scales.
The suppression of gravitational-wave echoes in nonlocal gravity does not originate from a selective suppression of ringdown frequencies. A crucial physical ingredient in the suppression mechanism is the invariant blueshift of the proper local frequency in the near-horizon region. For a mode of asymptotic frequency $\omega$, the locally measured frequency is given by
$\omega_{\rm loc}(r)=\frac{\omega}{\sqrt{f(r)}},$ (1)
where $f(r)=-g_{tt}(r)$. As $r\to r_{h}$ in a classical black hole geometry, $f(r)\to 0$ and $\omega_{\rm loc}\to\infty$. Even for a horizonless but highly redshifted compact object that mimics a black hole observationally, $f(r)$ becomes extremely small in the inner transition region, yielding $\omega_{\rm loc}\,\ell\gg 1$, where $\ell$ is the nonlocality length scale. It is this invariant blueshift, not a suppression of the observed gravitational-wave frequency, that activates the diffeomorphism-invariant entire-function regulator.
The suppression of ringdown echoes is a structural consequence of the smoothing of sharp reflective surfaces that would otherwise support repeated, phase-coherent backscattering. Echo formation requires the presence of nonanalytic features in the effective potential or boundary conditions, such as hard reflecting layers or discontinuous matching surfaces. The action of the entire-function regulator removes these features by replacing them with smooth, extended regions, thereby preventing the formation of a resonant echo cavity.
This mechanism operates independently of whether the underlying compact object possesses an event horizon. Both regular black holes with horizons and horizonless compact objects generated by nonlocal smearing exhibit the same qualitative behavior. The absence of sharply defined internal reflectors eliminates the conditions necessary for observable echoes. Although the nonlocal scale is taken to be Planckian, its role is not to suppress macroscopic frequencies, but to enforce analyticity and smoothness of the effective geometry. The resulting loss of phase coherence, rather than dissipative damping, is what removes the echo signal.
@@
The suppression of gravitational–wave echoes derived in this work operates entirely at the level of classical wave propagation on a fixed background geometry. The modified Regge–Wheeler–type equations [18, 19, 20], the smearing of effective potentials, and the exponential damping of reflection coefficients follow from classical integro–differential equations of motion. However, the specific nonlocal structure responsible for these effects, namely, analytic entire functions of the covariant d’Alembertian obeying Paley–Wiener bounds is not motivated by classical considerations alone. Such analyticity conditions are imposed to ensure unitarity, ghost freedom, and ultraviolet finiteness of the quantum gravitational theory. In this sense, quantum gravity is already implicitly encoded in the classical dynamics. The same analytic properties required for a consistent quantum theory manifest classically as a frequency–space filter that suppresses sharp reflections and eliminates the resonant cavity necessary for observable echoes. The resulting ringdown phenomenology reflects quantum–gravity consistency conditions, even in the absence of explicit quantum fluctuations or loop effects.
We show that these same analytic and diffeomorphism–invariant properties have direct and robust consequences for gravitational–wave ringdown phenomenology [13, 14, 15, 16, 17].
The purpose of this paper is to provide a clear and self-consistent account of echo suppression in analytic nonlocal gravity, emphasizing its geometric origin. By doing so, we aim to place the interpretation of gravitational-wave observations in nonlocal gravity on a firmer conceptual footing and to sharpen the distinction between nonlocal smoothing effects and phenomenological damping mechanisms.
@@ 2 Regular Static Spherically Symmetric Solution
and normalized so that the total mass is $M$:
@@
For the horizon phase structure only the ratio $GM/\ell$ matters.
@@
$x:=\frac{r}{\ell},\qquad\mu:=\frac{GM}{\ell},$ (15)
@@
3 Suppression of Ringdown Echoes in Covariant Nonlocal Quantum Gravity
@@
A static black-hole spacetime does not possess spatial translation invariance and therefore admits no global momentum-space representation of the form factor. The nonlocal regulator is defined covariantly as an entire function of the d’Alembertian, $\mathcal{F}\!\left(\Box\right)$, which is well-defined on any curved background and preserves diffeomorphism invariance. In a static geometry one retains time-translation symmetry and a conserved Killing frequency. In a stationary spacetime with timelike Killing vector $\xi^{\mu}$, the conserved Killing frequency is given by
$\omega\;\equiv\;-\,k_{\mu}\,\xi^{\mu}=k_{\mu}(\partial_{t})^{\mu},$ (23)
where $k^{\mu}$ is the wave four–vector. The locally measured proper frequency by an observer with four–velocity $u^{\mu}$ is given by
$\omega_{\mathrm{loc}}\;\equiv\;-\,k_{\mu}\,u^{\mu},$ (24)
where for static observers and integral curves of $\xi^{\mu}$, the normalized four–velocity is:
$u^{\mu}\;=\;\frac{\xi^{\mu}}{\sqrt{-\,\xi^{\alpha}\xi_{\alpha}}}.$ (25)
This allows for separation of variables and reduction to a one-dimensional Regge–Wheeler scattering problem. The action of $\mathcal{F}\!\left(\Box\right)$ is understood as a quasi-local smearing over a proper length scale $\ell$, rather than as a flat-space momentum filter. Echo suppression is consequently governed by adiabatic scattering controlled by the local parameter $\omega_{loc}\ell$, and does not rely on global translation invariance.
@@
where $K_{\Lambda_{G}}$ is a smooth kernel of width $\Delta r_{*}\sim\Lambda_{G}^{-1}$ in the local proper measure, determined by $\mathcal{F}$. Equation (27) states that the nonlocality smears the potential and modifies scattering. We define the smeared effective potential:
@@
The analytic nonlocality multiplicatively suppresses reflection through the factor $\widetilde{K}_{\Lambda_{G}}(2\omega)$. For the Gaussian filter:
@@
$R_{\rm ph,NL}(\omega)\;\approx\;\widetilde{K}_{\Lambda_{G}}(2\omega)\,R_{\rm ph,loc}(\omega).$ (32)
@@
where $T_{\mathrm{ph}}(\omega)$ denotes the complex transmission coefficient for a gravitational perturbation mode of frequency $\omega$ propagating through the photon-sphere and angular-momentum barrier in the Regge–Wheeler potential. $R_{\rm in}(\omega)$ is the inner reflection amplitude and $\Delta t$ is the round–trip time between the photon sphere and the inner scattering region expressed in tortoise coordinate. The transfer function is given by
@@
Combining (38) in the echo series (33), the $n$th echo carries a factor:
@@
$\Big|R_{\rm ph,NL}(\omega)\,R_{\rm in,NL}(\omega)\Big|^{n}\;\lesssim\;\exp\!\Big[-\,8n\,\omega^{2}/\Lambda_{G}^{2}\Big]\;\Big|R_{\rm ph,loc}(\omega)\,R_{\rm in,loc}(\omega)\Big|^{n},$ (41)
@@
Let us consider the role of the Paley–Wiener entire-function factor:
$\mathcal{F}(\omega)\equiv\exp\!\left(-\frac{\omega^{2}}{\Lambda_{G}^{2}}\right),$ (43)
which appears in nonlocal response functions and in the frequency-domain matching of Regge–Wheeler (RW) modes. For astrophysical ringdown, the relevant frequencies are set by the black-hole scale:
$\omega_{\rm QNM}\sim\mathcal{O}\!\left(\frac{1}{M}\right),\qquad f_{\rm QNM}\sim\mathcal{O}\!\left(\frac{1}{2\pi M}\right).$ (44)
For $M\sim 10\,M_{\odot}$ this corresponds to $f_{\rm QNM}\sim 10^{2}$$10^{3}\,$Hz. If $\Lambda_{G}$ is Planckian or even many orders below Planck, then $\omega_{\rm QNM}\ll\Lambda_{G}$ and we obtain:
$\mathcal{F}(\omega_{\rm QNM})=\exp\!\left(-\frac{\omega_{\rm QNM}^{2}}{\Lambda_{G}^{2}}\right)=1-\mathcal{O}\!\left(\frac{\omega_{\rm QNM}^{2}}{\Lambda_{G}^{2}}\right)\simeq 1.$ (45)
This explains why one should not expect any direct suppression of the black-hole frequency itself: the nonlocal scale is so high compared to $\omega_{\rm QNM}$ that (43) is essentially unity throughout the LIGO/Virgo/KAGRA band.
Echo phenomenology depends not on suppressing $\omega_{\rm QNM}$, but on the existence of a hard, highly reflective structure near the would-be horizon. In the RW problem, echoes arise only if there is an additional inner reflection coefficient $\mathcal{R}_{\rm in}(\omega)$ of magnitude close to unity, creating a cavity between the photon-sphere barrier and the inner surface. The observable echo train is controlled schematically by
$\tilde{Z}_{\rm echo}(\omega)\;\propto\;\frac{\mathcal{T}_{\rm ps}(\omega)\,\mathcal{R}_{\rm in}(\omega)\,e^{2i\omega L}}{1-\mathcal{R}_{\rm ps}(\omega)\,\mathcal{R}_{\rm in}(\omega)\,e^{2i\omega L}},$ (46)
where $L$ is the cavity length in tortoise coordinate, and $\mathcal{R}_{\rm ps},\mathcal{T}_{\rm ps}$ are the photon-sphere reflection and transmission amplitudes. The necessary condition for echoes is given by
$|\mathcal{R}_{\rm in}(\omega)|\not\ll 1\quad\text{over the ringdown band}.$ (47)
The nonlocal effect is that entire-function UV completion generically replaces a sharp distributional interaction localized at a surface by a smeared interaction of width $\ell\equiv\Lambda_{G}^{-1}$, so that would-be reflecting walls or abrupt transitions in effective medium properties are smoothed over a proper length $\sim\ell$. A convenient RW model is to represent the inner structure by an effective potential contribution localized near $r_{*}=r_{*0}$:
$V_{\rm in}(r_{*})=\lambda\,\delta(r_{*}-r_{*0})\qquad\longrightarrow\qquad V_{\rm in}^{\rm NL}(r_{*})=\lambda\,\frac{1}{\sqrt{\pi}\,\ell}\,\exp\!\left[-\frac{(r_{*}-r_{*0})^{2}}{\ell^{2}}\right],$ (48)
which is the action of an entire-function smearing map on a distributional source. For a monochromatic RW mode $\psi_{\omega}(r_{*})\sim e^{-i\omega t}$, the reflection amplitude off a localized perturbation is proportional to the Fourier transform of the profile. Consequently, the smearing in (48) yields the universal frequency dependence:
$\mathcal{R}_{\rm in}^{\rm NL}(\omega)\simeq\mathcal{R}_{\rm in}^{(0)}(\omega)\,\exp\!\left(-\frac{\omega^{2}\ell^{2}}{4}\right)=\mathcal{R}_{\rm in}^{(0)}(\omega)\,\exp\!\left(-\frac{\omega^{2}}{4\Lambda_{G}^{2}}\right),$ (49)
up to model-dependent phases and prefactors, where $\mathcal{R}_{\rm in}^{(0)}$ is the unsmeared local inner reflectivity. Equation (49) shows the correct physical separation of roles. The nonlocal factor does not suppress the ringdown frequency $\omega_{\rm QNM}$ itself as demonstrate in (45). Instead, it suppresses the ability of a sharp surface to reflect RW modes, by exponentially damping of the proper local frequency $\omega_{loc}$ components associated with an abrupt impedance mismatch. In other words, nonlocality eliminates the hard reflecting boundary condition by turning it into a smooth kernel.
4 Regge–Wheeler Matching and Vanishing of the Inner Reflection Coefficient
The frequency domain, $\Psi_{\omega\ell}(r_{*})$ is governed by the Regge–Wheeler equation:
$\left[\frac{d^{2}}{dr_{*}^{2}}+\omega^{2}-V_{\rm eff}(r)\right]\Psi_{\omega\ell}(r_{*})=0,\qquad\frac{dr_{*}}{dr}=f^{-1}(r),$ (50)
where $f(r)$ is the background lapse function and $V_{\rm eff}(r)$ is the standard axial or polar effective potential. In models that admit gravitational-wave echoes, an effective inner boundary condition is imposed at some $r_{*}=r_{*}^{(0)}$:
$\Psi_{\omega\ell}(r_{*})\;\xrightarrow[r_{*}\to r_{*}^{(0)}]{}e^{-i\omega r_{*}}+R_{\rm in}(\omega)\,e^{+i\omega r_{*}},$ (51)
where $R_{\rm in}(\omega)$ represents the inner reflection coefficient. A nonvanishing $R_{\rm in}$ signals the presence of a sharp reflective structure capable of generating late-time echoes.
In analytic nonlocal gravity, the dynamics is modified by an entire-function form factor involving the covariant d’Alembertian. At the level of the one-dimensional scattering problem, this implies that sharp structures in the effective potential are replaced by a smeared, analytic profile,
$V_{\rm eff}(r_{*})\;\longrightarrow\;V_{\rm eff}^{\rm NL}(r_{*})\equiv e^{\Box/\Lambda_{G}^{2}}\,V_{\rm eff}(r_{*}).$ (52)
In a locally flat approximation across the narrow inner transition region, the action of the nonlocal operator may be written as a convolution:
$V_{\rm eff}^{\rm NL}(r_{*})=\int_{-\infty}^{+\infty}ds\;\frac{1}{\sqrt{4\pi}\,\ell}\,e^{-s^{2}/(4\ell^{2})}\,V_{\rm eff}(r_{*}-s),\qquad\ell\equiv\Lambda_{G}^{-1}.$ (53)
Equation (53) shows explicitly that any nonanalytic feature—such as a thin shell, hard surface, or discontinuous matching layer—is replaced by a smooth transition region of width $\mathcal{O}(\ell)$.
To extract the inner reflection coefficient generated by the smooth transition, we rewrite Eq. (50) in Lippmann–Schwinger form:
$\Psi_{\omega\ell}(r_{*})=e^{-i\omega r_{*}}+\int_{-\infty}^{+\infty}dr_{*}^{\prime}\;G_{\omega}(r_{*}-r_{*}^{\prime})\,V_{\rm eff}^{\rm NL}(r_{*}^{\prime})\,\Psi_{\omega\ell}(r_{*}^{\prime}),$ (54)
where the one-dimensional Green function is given by
$G_{\omega}(x)=\frac{i}{2\omega}\,e^{i\omega|x|}.$ (55)
In the Born approximation, sufficient to determine the scaling of the reflected amplitude, one substitutes $\Psi_{\omega\ell}\to e^{-i\omega r_{*}}$ in the integral. The coefficient of the outgoing wave $e^{+i\omega r_{*}}$ on the inner side then yields:
$R_{\rm in}(\omega)=\frac{1}{2i\omega}\int_{-\infty}^{+\infty}dr_{*}\;e^{2i\omega r_{*}}\,V_{\rm eff}^{\rm NL}(r_{*}).$ (56)
Thus, $R_{\rm in}(\omega)$ is proportional to the Fourier transform of the inner transition profile evaluated at wavenumber $2\omega$. Reflection therefore directly probes the sharpness of the effective potential.
To make this explicit, consider a localized inner structure whose nonlocal image is well approximated by an analytic profile:
$V_{\rm eff}^{\rm NL}(r_{*})\simeq V_{0}\,\exp\!\left[-\frac{(r_{*}-r_{*}^{(0)})^{2}}{\ell^{2}}\right].$ (57)
Substitution into Eq. (56) yields:
$R_{\rm in}(\omega)=\frac{V_{0}}{2i\omega}\,\sqrt{\pi}\,\ell\,e^{2i\omega r_{*}^{(0)}}\,\exp\!\left[-\omega^{2}\ell^{2}\right].$ (58)
The exponential factor in Eq. (58) arises solely from the analyticity of the smeared potential and represents suppression of the reflection coefficient, not suppression of the physical ringdown frequencies.
Although $\ell$ is Planckian, the relevant quantity controlling reflection is the local proper frequency. Near a would-be horizon:
$\omega_{\rm loc}(r)=\frac{\omega}{\sqrt{f(r)}}\;\xrightarrow[r\to r_{h}$$\infty$. (59)
Expressed in terms of proper distance across the smooth transition region, Eq. (58) implies
$|R_{\rm in}(\omega)|\;\sim\;\exp\!\left[-\omega_{\rm loc}^{2}\,\ell^{2}\right]\;\longrightarrow\;0.$ (60)
Therefore, the inner reflection coefficient vanishes because the entire-function regulator removes sharp reflective structures from the effective geometry. The absence of echoes follows from the loss of phase-coherent inner reflection, not from any suppression or filtering of black-hole quasinormal-mode frequencies.
The quantity that controls reflection from the inner transition region is not the asymptotic frequency $\omega$ measured by a distant observer, but the local proper frequency measured by a physical observer at radius $r$. For a static, spherically symmetric background:
$ds^{2}=-f(r)\,dt^{2}+f^{-1}(r)\,dr^{2}+r^{2}d\Omega^{2},$ (61)
a perturbative mode with time dependence $\Psi\sim e^{-i\omega t}$ has locally measured frequency:
$\omega_{\rm loc}(r)=\frac{\omega}{\sqrt{f(r)}}.$ (62)
This is the standard gravitational blueshift of frequencies toward regions of large redshift.
The relevance of $\omega_{\rm loc}$ follows from the fact that mode conversion between ingoing and outgoing solutions is governed by the ratio of the local oscillation scale to the length scale over which the effective potential varies. In analytic nonlocal gravity, the entire-function regulator smooths any would-be inner reflective structure over a proper length $\ell\equiv\Lambda_{G}^{-1}$, so that the dimensionless parameter controlling reflection is $\omega_{\rm loc}\ell$, rather than $\omega\ell$.
For a black hole with an event horizon at $r=r_{h}$, one has $f(r_{h})=0$, and it follows:
$\omega_{\rm loc}(r)=\frac{\omega}{\sqrt{f(r)}}\;\xrightarrow[r\to r_{h}$$\infty$. (63)
As a result, even gravitational waves with astrophysically small asymptotic frequencies acquire arbitrarily large local proper frequencies in the vicinity of the horizon. Since the inner transition region is smoothed on a scale $\ell$, this implies:
$\omega_{\rm loc}\,\ell\gg 1,$ (64)
and therefore an exponential suppression of the inner reflection coefficient.
The same mechanism operates in regular, horizonless compact objects generated by nonlocal smearing. Although $f(r)$ does not vanish, such geometries typically contain a deep redshift well, characterized by a minimum value $f_{\rm min}\ll 1$ in the inner transition region. The corresponding maximal local frequency is given by
$\omega_{\rm loc}^{\rm max}=\frac{\omega}{\sqrt{f_{\rm min}}},$ (65)
which can be enormously larger than $\omega$ even for moderate redshifts. Consequently, for Planckian smoothing lengths $\ell\sim\ell_{\rm Pl}$ one again finds:
$\omega_{\rm loc}^{\rm max}\,\ell\gg 1,$ (66)
ensuring that the inner transition region is traversed adiabatically and that ingoing and outgoing modes do not mix.
In both horizon and horizonless cases, the limit $\omega_{\rm loc}\ell\gg 1$ implies that the effective geometry is impedance matched across the inner transition region, yielding:
$R_{\rm in}(\omega)\;\to\;0.$ (67)
The suppression of gravitational-wave echoes therefore arises from the loss of phase-coherent inner reflection enforced by nonlocal smoothing, rather than from any suppression or filtering of the physical ringdown frequencies themselves.
This is why the same UV completion can leave classical Kerr-Schwarzschild ringdown essentially unchanged, while still predicting no observable echoes from regular and horizonless compact objects. Echoes require $|\mathcal{R}_{\rm in}|\sim 1$, whereas entire-function smearing drives $|\mathcal{R}_{\rm in}^{\rm NL}(\omega)|\ll 1$ across the relevant band whenever the would-be surface is not fundamentally sharp. Operationally, the nonlocal theory replaces a Dirichlet-Neumann-type inner boundary condition at $r_{*0}$ by a nonlocal relation of the form:
$\psi_{\omega}(r_{*0})\;\;\longrightarrow\;\;\int_{-\infty}^{+\infty}\!{\rm d}r^{\prime}_{*}\;K_{\ell}(r_{*0}-r^{\prime}_{*})\,\psi_{\omega}(r^{\prime}_{*}),\qquad K_{\ell}(x)=\frac{1}{\sqrt{\pi}\,\ell}\,e^{-x^{2}/\ell^{2}},$ (68)
which removes the hard wall responsible for large reflection in the local horizonless models.
The cavity phase $e^{2i\omega L}$ in (46) is unaffected at leading order when $\omega\ll\Lambda_{G}$; the suppression mechanism acts through $\mathcal{R}_{\rm in}(\omega)$, not by shifting the characteristic black-hole frequencies. The nonlocality does not suppress the black-hole ringdown frequency, but it suppresses the reflectivity of any near-horizon or horizonless surface by smoothing it over the nonlocal length $\ell$, thereby eliminating observable echoes.
The suppression of echoes in the present framework follows from two independent but physically linked ingredients: (i) smoothing of the inner transition region by the nonlocal entire-function regulator over a proper length scale $\ell$, and (ii) the extreme gravitational redshift required for black-hole mimicry.
Consider a horizonless compact object whose surface or inner transition region lies at:
$R=2M(1+\epsilon),\qquad 0<\epsilon\ll 1.$ (69)
The exterior metric function satisfies:
$f(R)=1-\frac{2M}{R}=\frac{\epsilon}{1+\epsilon}\simeq\epsilon,$ (70)
so that the gravitational redshift at the surface is given by
$1+z=\frac{1}{\sqrt{f(R)}}\simeq\epsilon^{-1/2}.$ (71)
An object that is observationally indistinguishable from a classical black hole must have $\epsilon\ll 1$ in order to suppress detectable surface emission and luminosity from infalling matter. Extreme redshift is therefore not an additional assumption but a necessary condition for black-hole mimicry.
A perturbation mode of Killing frequency $\omega$ measured at infinity is locally measured in the inner transition region with proper frequency:
$\omega_{\rm loc}(R)=\frac{\omega}{\sqrt{f(R)}}\simeq\frac{\omega}{\sqrt{\epsilon}}.$ (72)
Even for modest astrophysical ringdown frequencies $\omega$, the locally measured frequency becomes parametrically large as $\epsilon\to 0$.
The reflection coefficient from a smooth transition region of proper thickness $\ell$ depends on the dimensionless adiabaticity parameter:
$x\equiv\omega_{\rm loc}\,\ell=\frac{\omega\ell}{\sqrt{\epsilon}}.$ (73)
When $x\gg 1$, the mode probes the transition region in the adiabatic regime and the inner reflection coefficient is strongly suppressed:
$R_{\rm in}(\omega)\to 0.$ (74)
Ultra-compact horizonless configurations that genuinely mimic black holes have $\epsilon\ll 1$. The large gravitational blueshift ensures $\omega_{\rm loc}\ell\gg 1$ even if $\omega\ell\ll 1$ at infinity. Echo suppression therefore arises from the combination of extreme redshift and nonlocal smoothing of sharp reflective structures, rather than from any suppression of the asymptotic ringdown frequency itself.